### v1

# Automatic Fact-Checking with Frame-Semantics

Abstract
We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims, addressing the challenges posed by misinformation in today’s information ecosystem. To support this approach, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data. This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames and data sources from the Organisation for Economic Co-operation and Development (OECD). Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval, indicating a meaningful advancement in automatic fact-checking capabilities. Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future research.
Automatic Fact-Checking with Frame-Semantics
Jacob Devasier, Rishabh Mediratta, Akshith Putta, Chengkai Li University of Texas at Arlington cli@uta.edu
1 Introduction
The proliferation of misinformation presents a significant challenge to the modern information ecosystem. Mitigating the spread of fake news has become a critical concern for researchers, policymakers, and technology developers alike. In response, automatic fact-checking has emerged as an important research area, with the goal of reducing the burden on human fact-checkers by automating key steps in the fact-checking pipeline.
Many previous studies on automatic fact-checking have utilized unstructured data, such as fact-checks from trustworthy sources Wang (2017), to perform claim matching Shaar et al. (2020) and use large language models to provide verdicts and explanations Cheung and Lam (2023b); Singhal et al. (2024a); Khaliq et al. (2024). Some works have also studied structured data, such as tabular data from Wikipedia Chen et al. (2020); Aly et al. (2021) or scientific documents Wang et al. (2021); Akhtar et al. (2022). These works have been studied on both simple claims extracted from Wikipedia Bouziane et al. (2021) and real-world claims Wang et al. (2021); Akhtar et al. (2022).
Previous approaches Ye et al. (2023); Karagiannis et al. (2020); Jo et al. (2019) have found success with using SQL to query Wikipedia tables, which have an average of 13 rows per table Chen et al. (2020). color=green]The last sentence “due to a lack of …” is a bit confusing . I suppose we are talking about Wikipedia tables only. If so, then the notion of “datasets” is moot, because there is only one dataset that’s applicable—all tables in Wikipedia. By “due to a lack of ”, did you intend to say that most tables there are small? — the point I was making is that there are no existing large-scale datasets for structured automatic fact checking, which is why no past works have addressed it Due to a lack of existing datasets, no previous work has studied automatic fact-checking on real-world claims using large-scale structured data. To overcome the limitations of previous methods, we introduce a novel pilot dataset of real-world claims extracted from PolitiFact  [^1] https://www.politifact.com/ annotated for large-scale structured data.
Our proposed pilot dataset serves as the foundation for two case studies focused on leveraging frame semantics for automatic fact-checking. Frame semantics Ruppenhofer et al. (2016); Baker et al. (1998) is a linguistic framework that explores how language encodes meaning through structured representations called frames. These frames capture the essential elements and relationships in various situations, enhancing the understanding of context and intent behind statements.
In the first case study, we utilize the Vote semantic frame Arslan et al. (2020), which represents how an Agent (e.g., a politician) interacts with a particular Issue (e.g., a proposed bill) through voting. Each frame includes specific components known as frame elements, which provide additional context about the roles and relationships involved, such as the voting member, the bill at stake, and the outcome of the vote. color=cyan]This sentence “Each frame includes …” is a little problematic. The previous sentence already discussed Agent and Issue, which are frame elements. And this sentence is a bit redundant with the previous paragraph. Some re-ordering, reorganization may help. This case study includes 79 claims along with their corresponding congressional members and bills, with an average of 4,230 votes per congress member extracted from official U.S. voting records.color=green]you mentioned the frame side of the framework, but not the database side of it. Particularly, the 4,230 votes per member come from the voting record database. Without discussing the database and its role in this, a reviewer may get confused by the mentioning of the number.
The second case study broadens the scope by exploring a variety of additional semantic frames and a diverse collection of datasets from the Organisation for Economic Co-operation and Development (OECD), which contain over 400 tables with an average of 596,000 rows and 13 columns. color=green]should we say “OECD-related claims” instead? Also, “For OECD claims, we annotated 68 claims” is strange. Couldn’t we just say “We annotated 68 OECD-related claims”? We annotated 68 OECD-related claims along with their corresponding frames and mapped each claim to the OECD statistics used to fact-check them. color=green]The notion of “evoke” hasn’t been mentioned. color=green]Similar to the last paragraph, you haven’t discussed the role of the OECD data in fact-checking. These case studies help to identify the challenges associated with automated fact-checking using large-scale structured data.
Our findings demonstrate the value of frame semantics in automatic fact-checking by enabling a structured and explainable understanding of claims. For instance, on voting claims, we found that using frame elements extracted from claims, instead of the entire claim itself, improved evidence retrieval, leading to a +2.1-point increase in recall@10 for voting claims. Similarly, on OECD claims, we saw a +7.3-point increase in recall@5 when frame elements were used to identify relevant OECD tables. These improvements illustrate how frame-semantics color=cyan]earlier you only said “frame semantics”, not parsing. can enhance retrieval by guiding systems toward relevant structured data.
To understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and found a heavy skew towards a few specific frames, including Taking_sides, Speech, Change_position_on_a_scale. These insights guided the selection of frames for our case studies and enable future works to study high-impact frames.
To summarize, our contributions are as follows:
  • We proposed a novel paradigm for automatic fact-checking using frame-semantics.
  • We developed a pilot dataset for automatic fact-checking of real-world claims using large-scale structured data. [^2] https://github.com/anon-naacl25/
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, allowing researchers to target high-impact frames for future studies.
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^3] https://anon-naacl25.github.io/
2 Related Works and Background
Recent advances in automatic fact-checking have been largely driven by the integration of large language models (LLMs) and retrieval-augmented generation (RAG) pipelines. Wang et al. (2024) proposed a unified framework for LLM-based systems which utilizes an internal mechanism to determine which of a selection of LLM-base models should be used to fact-check a particular claim. RAGAR Khaliq et al. (2024) enhances fact-checking by using multimodal inputs and iterative reasoning. Similarly, FactLLaMA Cheung and Lam (2023a) combines pre-trained LLaMA models with external evidence retrieval to verify claims. LLM-Augmenter Peng et al. (2023) integrates external knowledge sources and providing feedback to improve model accuracy. Singhal et al. (2024b) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
For structured data, many approaches have been studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2022) for fact verification. Dater Ye et al. (2023) is a detailed fact verification system which decomposes claims into sub-questions using LLMs to simplify the fact-checking process and is the best-performing model on the TabFact Chen et al. (2020) dataset. In this work, we do not focus on fact verification; however, we choose to utilize LLM-based methods for our fact verification step as they do not require retraining the model when new data is used. color=green]I felt this statement undermines the significance of our work. It focuses on the usage of LLM. But the more important essence of our work is the usage of frames.
2.1 End-to-end Fact-Checking Systems
ClaimBuster Hassan et al. (2017) describes the process of automatic fact-checking as having several core components including a claim matching component, where claims are matched with previous fact checks if they exist, and a knowledge base lookup, which can directly answer questions using internal knowledge, e.g., Wolfram|Alpha. Automatic fact-checking systems, like this work, aim to fulfill the role of the knowledge base. color=cyan]Is that internal or external? More recent studies Guo et al. (2022) have decomposed the process into claim detection, evidence retrieval, and verdict prediction and justification. color=green]Looks strange to have a subsection for a short paragraph. Its connection to the whole section and to the paper’s topic is not clearly explained. — paragraph header looks a bit weirder in my opinion. should we just remove this subsection?
Figure 1: An example of our proposed paradigm. First, the frame and frame elements (FEs) are extracted using a frame-semantic parser (FSP). Then, FEs are mapped to a table which can then be queried to gather evidence for the claim. Finally, the evidence and claim are passed into a fact verification model to check the truthfulness of the claim.
3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment. Figure 1 provides an example of how the system processes the statement “Japan’s life expectancy is the 2nd best in Asia” to illustrate this process.
color=green]Need some better re-organization. the first half of this section has several segments, each with a bold-face title. Then we suddenly start to have subsections for the 2nd half, where each subsection is just a paragraph.
3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing enables the extraction of structured, predefined representations from claims. By identifying frames and their associated frame elements (FEs), the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification. For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.
Frame semantics not only facilitate claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables, other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.
3.2 Data Collection and Evidence Retrieval
The availability of trustworthy data is essential for fact-checking. Rather than relying on search engines that can return inconsistent results or be vulnerable to fake news injections Huynh and Hardouin (2023); Horne et al. (2019), our system stores ground-truth data in an internal database. Using data from reliable sources like the OECD and U.S. Congress ensures the integrity of the evidence.
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy. Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by adapting methods like retrieval-augmented generation Lewis et al. (2020) using text chunking techniques to query large text corpora instead of a database.
3.3 Claim-Evidence Alignment
The final step involves aligning the retrieved evidence with the original claim. Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context. By only using specific spans of the claim, the system can align the claim and evidence more precisely. This approach also enhances explainability, as the system can point to the exact text used in the query.
Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim. For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best. While this claim is not exactly true, the spirit of the claim is true, i.e., that the life expectancy is extremely high in Japan, and thus the model would predict that it is mostly true.color=yellow]This is an intriguing example. But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it? — this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
While fine-tuned approaches have shown slightly better performance on table-based fact-checking Ye et al. (2023), we employ an LLM-based approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions. Because this step is not a key focus of our work, we leave fact verification model-agnostic to allow future improvements. Thus, we can replace the current LLM with any fact verification model which takes structured evidence and a claim to perform fact verification.
4 Fact-Checking Case Studies
In this section we provide two case studies which utilize our proposed paradigm.
4.1 Voting Records
PolitiFact is a leading fact-checking organization that assesses the accuracy of public statements using verifiable evidence. Voting records  [^4] https://www.politifact.com/voting-record/ are a substantial portion of PolitiFact’s fact-checks. In this case study, we focus on automatically fact-checking voting-related claims using the Vote frame and official U.S. congressional records. Specifically, we target the Agent and Issue Frame Elements (FEs), which refer to the voting entity and the topic of the vote, respectively.
Datasets.
We compiled a dataset of official U.S. congressional voting records from the Congress Github Repository. [^5] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.
In their fact-checking process, PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims. In the context of the Vote frame, claims often reference specific congressional bills. To evaluate our system, we constructed an evaluation dataset by extracting all PolitiFact fact-checks made before April 2022 that involve claims related to the Vote frame and reference at least one congressional bill. After manual verification, we collected 79 fact-checks, along with their corresponding bills, to form the evaluation set.
Congress Member Identification.
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use SQL queries to match congress members whose names are similar to the words in the Agent FE. In cases of name ambiguity, e.g., when two members share the same name, we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Although not exhaustive, these lists should cover many common cases. Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz. We handle this by utilizing a list of congress members’ preferred names, supplemented by common alternatives  [^6] https://github.com/carltonnorthern/nicknames when necessary.
Bill Matching.
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing.
Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, which allows us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
Verifying alignment of claim to bill.
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (FE) (for or against). Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”. Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
Claim Verification.
Finally, our system integrates the alignments between each bill and claim to perform fact verification over all of the retrieved evidence. We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix D.4. color=green]Not clear to me what “each individual claim-bill alignement” means.
Figure 2: System design of OECD case study. The frame-semantic parser (orange) facilitates the claim understanding process, table matching and querying (blue) facilitate the evidence retrieval process, and the claim verification step (green) facilitates the fact verification.
4.2 OECD-Related Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics and serves as a strong focal point to explore a wider range of semantic frames. As a visual aid, we include a system overview in Figure 2.
Dataset.
For the OECD Case Study, we collect all of the data tables available on the OECD Data Explorer.  [^7] https://data-explorer.oecd.org/ We construct a SQLite database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totalling 4.1 billion cells.
Relevant Table Identification.
To identify relevant tables for data extraction, we apply semantic similarity models to select the five most similar tables based on the extracted frame elements (FEs). The query consists of one or more FEs derived from frame-semantic parsing, while the target is a specific table. We encode each table’s name and description as text for this purpose. Given the brevity of our queries and the length of the table descriptions, we utilize asymmetric semantic similarity models to ensure precise matches. To minimize the risk of overlooking relevant data, we retrieve the top five most similar tables.color=green]Is “asymmetric semantic similarity” a research field, a technique, or a particular model (maybe in GPT suite or library)? Need to be clear on this and provide link/reference. — a type of model determined by how it is trained, i.e., short query vs long document
Querying Relevant Tables.
Querying the relevant tables presents challenges, as the system must recognize how values in the claim are represented in the database. color=cyan]We said “presents challenges” quite many times in the system. You might want to consider add some variety to the wording. In cases where direct mappings are available (e.g., in the Vote frame case study), we map FEs to specific tables. However, in the OECD case study, where no direct mapping exists between FEs and table columns, the system must predict this mapping. To handle these cases, we employ LLM-based code generation, allowing for the execution of multiple simple queries using programmed logic, rather than relying on complex text-to-SQL generation.color=cyan]Should we include some text-to-SQL references?
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2023), which studied different methods of representing SQL tables for text-to-SQL generation.color=green]What is Gao et al. 2023 about? on the same problem or something else. YOu might want to make this clear. Alongside the schema, we include representative example values from the database for each column. color=green]“Columns …. include all values” doesn’t seem correct. — referring to how they are encoded into text Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values. For columns with a moderate number of unique values, we use a semantic similarity model to select the top ten most relevant values based on the claim. Given the large number of cells generated by this process, we use an LLM to refine the query and filter out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompt for this LLM can be found in Appendix D.3. color=green]Is this semantic similarity model the same as the aymmetric one mentioned in the last section? If yes, make it clear; if not, need reference/link, based on the same concern mentioned earlier. — each model is different because they perform differently on the data, the selection is discussed in the experiments color=green]Are we missing some details here about how we used LLM? Are there prompts?
Fact Verification
In the final step, we used an LLM-based fact verification model. We represent the extracted evidence as a list of tab-separated values based on the data retrieved in the previous step. The model is instructed using the prompt in Appendix D.4 and outputs one of five verdicts: false, mostly false, half-true, mostly true, or true.color=cyan]If time allows, explain the connection to PolitiFact’s Truth-o-meter.
5 Experiments
5.1 Datasets
Fact-checking Frames.
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in this dataset along with the newly-defined frames. These frames can be found in Appendix 6.
PolitiFact Fact-checks
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles collected as of April 2022. Each article contains a claim, a detailed fact-check, a verdict, and a list of sources. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we identified 79 fact-checks that cite congress.gov and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset.
color=cyan]PolitiFact fact-checks and PolitiFact fact-check articles. Mention very close to each other. color=green]the claim cites congress.gov? color=green]“bill cited was referenced by the claim”. cited where? also the claim would reference a bill?
Additionally, we collected 68 articles that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the statistics which can be used to fact-check it. These two sets of claim-evidence pairs—the voting-related claims and the OECD-related claims—serve as evaluation sets for our case studies. color=green]these articles cite the website? color=green]“To collect … we collect” color=green]Which two sets?
Model Frames Frame Acc FE Acc
Random Vote 0.488 0.254
GPT-4o-mini Vote 0.974 0.618
Vote FSP Vote 0.990 0.889
Random OECD 0.602 0.000
GPT-4o-mini OECD 0.537 0.372
GPT-4o-mini* OECD 0.713 0.461
OECD FSP OECD 0.742 0.873
Table 1: Performance of frame-semantic parsing model on fact-checking frames alongside simple baseline.
Frame Samples (%)
Taking_sides 7,152 (34.0%)
Speech 6,010 (28.6%)
Change_position_on_a_scale 5,547 (26.4%)
Comparing_two_entities 5,530 (26.3%)
Cause_change_of_position_on_a_scale 4,675 (22.2%)
Vote 3,229 (15.4%)
Comparing_at_two_different_points_in_time 2,436 (11.6%)
Conditional_occurrence 2,355 (11.2%)
Creating 2,194 (10.4%)
Occupy_rank 1,106 (5.3%)
Oppose_and_support_consistency 1,010 (4.8%)
Recurrent_action_in_Frequency 935 (4.4%)
Ratio 932 (4.4%)
Capability 869 (4.1%)
Occupy_rank_via_superlatives 767 (3.6%)
Uniqueness_of_trait 497 (2.4%)
Occupy_rank_via_ordinal_numbers 329 (1.6%)
Recurring_action 187 (0.9%)
None 12 (0.1%)
Table 2: Distribution of semantic frames in PolitiFact.
5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combine the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unify frame and frame element identification into a single model. This model is then trained only on the fact checking frames introduced in Arslan et al. (2020).
To evaluate the performance of our model for the case studies, we compare it with a GPT-4o-mini model using OpenAI’s structured generation [^8] At the time of submission, GPT-4o-mini is the only model which supports structured generation.. We evaluate each model using accuracy on frame identification and exact match accuracy on FE identification in Table 1. We define a separate prompt for each case study in Appendix D.1.
5.3 Claim Coverage
One concern which may arise is regarding the applicability of our system for assisting in the fact checking process. To understand this, we aim to identify the coverage of a frame-based approach towards automatically fact-checking claims. Due to the lack of annotations for all of the fact-checking frames in Arslan et al. (2020), we utilize a zero-shot GPT-4o-mini model to identify the frames evoked in the claim of each PolitiFact fact-check (Table 2). Claims which do not evoke a frame studied in this work are categorize as none. The prompt used for this model can be found in Appendix D.1.
Model Data R@K
BM25 OECD 0.323
distilbert-tas-b OECD 0.505
text-embedding-3-large OECD 0.610
roberta-base-v2 OECD 0.726
text-embedding-3-large Vote 0.032
stella_en_1.5B_v5 Vote 0.144
distilbert-tas-b Vote 0.165
Table 3: Performance comparison of different semantic similarity models on extracted frame elements. K=10 for Vote and K=5 for OECD.
Model Query Data R@K
RoBERTa (v2) Full claim OECD 0.653
RoBERTa (v2) FE OECD 0.726
Max Possible - OECD 0.910
distilbert-tas-b Full claim Vote 0.143
distilbert-tas-b Issue FE Vote 0.165
Max Possible - Vote 0.568
Table 4: Performance of different methods of representing the query in the OECD case study’s table matching. OECD and Vote claims use Recall@5/10, respectively.
5.4 Evidence Retrieval
We study the capability of our system to retrieve the relevant information which can be used to fact-check the claim. To do this, we found the percentage of bills and statistics available in our database for voting and OECD-related claims, respectively. This serves as a performance ceiling on our collected claims as our system would be unable to fact-check claims which do not have available data. These values are found in Table 4 (Max Possible).
For semantic similarity, we compare commonly used embedding models (Table 3), namely RoBERTa Liu et al. (2019) and DistilBERT-TAS-B Hofstätter et al. (2021). Additionally, we compare with OpenAI’s text-embedding-3-large model and provide a baseline BM25 approach. We also experiment with different approaches of representing the query used to find relevant data. In Table 4, we compare the performance of using the entire claim with only the relevant FE. To evaluate the performance of these queries, we calculate the recall@K, where K=5 for OECD claims and K=10 for Vote claims. The relevant FEs used for each frame can be found in Appendix A.
Model Dataset Accuracy
GPT-4o Naive Vote 0.044
Our w/ Irrelevant Vote 0.076
Our w/o Irrelevant Vote 0.207
Our w/ Irrelevant ± Vote 0.253
GPT-4o Naive ± Vote 0.324
Our w/o Irrelevant ± Vote 0.690
GPT-4o Naive OECD 0.073
Our w/ Irrelevant OECD 0.214
Our w/o Irrelevant OECD 0.429
GPT-4o Naive ± OECD 0.537
Our w/ Irrelevant ± OECD 0.607
Our w/o Irrelevant ± OECD 0.821
Table 5: Fact verification performance on different case studies. ± indicates the off-by-one performance. Bolded values indicate the best-performing settings overall and underlined values are only considering exact matches.
5.5 Fact Verification
Finally, we study the ability of our fact verification system to fact-check claims using structured data by comparing the verdicts presented in each PolitiFact fact-check to the model’s predicted verdict. We have replaced PolitiFact’s “pants on fire” verdicts with a false verdict. To determine whether the LLM is relying on internal knowledge or if it is using the provided data, we evaluate the performance of GPT-4o without any data. We measure accuracy as the primary evaluation metric and also calculate off-by-one accuracy, where, for example, a prediction of “mostly false” is considered correct if the label is “false.” The results of these experiments are summarized in Table 5.
6 Results
6.1 Frame-Semantic Parsing
We evaluated the frame-semantic parsing model on fact-checking frames introduced by Arslan et al. (2020), comparing it against GPT-4o-mini due to its structured output capability. Table 1 shows the performance on both the Vote and OECD frames. We found that the GPT-4o-mini model tends to over-predict the number of frames in a sentence. When we only use the first predicted frame, the performance is increased significantly. We saw similar cases with frame element predictions where GPT-4o-mini tends to predict more frame elements than actually exist in the input. We include examples of each of these cases in Appendix C.2.
6.2 PolitiFact Survey
Similar to Section 6.1, we found that GPT-4o-mini tends to predict more frames than there may actually be. We believe this is likely due to the annotations from Arslan et al. (2020) being sampled from PolitiFact and likely having a similar structure. We found the annotated samples evoked 1.1 frames on average while GPT-4o-mini predicts an average of 2.1 frames. Another possible explanation is that the model is able to identify frames for lexical units– words which evoke frames–which are not defined.
One indication of this is on the Vote frame, which only has a lexical unit defined for vote.v; however, it is possible to evoke the Vote frame using other words. For example, the statement “I passed a bill” implies that a vote of affirmation was cast on that bill. We found only 1,198 (5.7%) PolitiFact fact-checks which mention “vote” while the model predicts that almost 3 times as many evoke the Vote frame. While we found many claims which evoke a Vote frame without an explicit mention of the word “vote”, further analysis is needed to get a more accurate understanding of the true distribution of frames evoked by factual claims.
6.3 Evidence Retrieval
We found that embedding models’ performance is task-specific and there is no single-best embedding model for every task. For example, RoBERTa performed the best for finding the right OECD table while DistilBERT-TAS-B was better for matching claims to their corresponding bills. We also found that, in general, using the claim-specific frame element performs better than using the entire claim as a query, as shown in Table 4.
For voting-related claims, we found that identifying the bill referenced in a particular claim is very difficult. Our best performing model, DistilBERT-TAS-B, was only able to obtain a Recall@10 of 0.165. For OECD-related claims, identifying the table to use to fact-check the given claim was much easier, with the best model, RoBERTa, achieving a recall@5 of 0.726. This is likely because the similarity between the relevant FE and the table names were better-represented by semantic similarity models. This indicates a potential area of significant improvement in the methodology and representation of bill texts according to their topics and implications, in addition to the text itself.
6.4 Fact Verification
On OECD claims, 62% of the generated queries were able to retrieve relevant data while only 36% of the voting claims were able to retrieve relevant bills. These values put an upper bound on the capabilities of our fact verification model as without evidence, the model will not be able to provide useful fact-checks. In general, we found that language models tend not to predict a claim to be true based on the given data, yet they have no issue with predicting a claim to be false, and often do so. We believe this is analogous to proofs, it is quite easy to prove by contradiction, if you find a single vote or country statistic which contradicts the claim, while it is much harder to assert that a claim is true without looking at all of the data.
We found that on Vote claims, a naive GPT-4o correctly predicted the truthfulness of the factual claim with an accuracy of 0.286; however, when voting data is added, the performance drops to 0.076. This suggests that, while there may be some internal knowledge on certain people and bills, it is unlikely that the model is relying on it when given specific voting records. This also indicates that the prompt given to the model is sufficient in preventing the LLM from relying on internal knowledge.
When excluding claims for which data was not found, we observed absolute performance increases of +13.1% and +21.5% for Vote and OECD claims, respectively. We also found that, while the exact-match accuracy of the system is low, there is a strong correlation between predictions and labels, as indicated by the off-by-one accuracy in Table 5. We include example outputs for the explanations of these models in Appendix C.1.
7 Conclusion
This research contributes to the evolving field of automatic fact-checking by proposing a novel paradigm that incorporates frame semantics to enhance the evidence retrieval process and enable fine-grained control of downstream tasks. By developing and presenting a pilot dataset specifically annotated for real-world claims, this work helps to address the underexplored challenge of fact-checking using massive, structured data sources.
Our findings demonstrate the promise of frame-semantic parsing in improving evidence retrieval for automatic fact-checking. In both case studies, we observed that using frame elements extracted from claims instead of entire claims led to substantial improvements in recall metrics, highlighting the utility of frame-based understanding in identifying relevant evidence. These results suggest that semantic frames not only offer a structured and explainable approach to parsing claims but also guide retrieval models more effectively toward the correct information within large datasets.
Looking ahead, the integration of frame semantics into fact-checking systems holds promise for further advancements in both accuracy and interpretability. Future work could explore broader applications of this approach, including expanding the dataset, refining the models, and incorporating additional semantic frames to cover more claims.
Limitations
Despite the advancements made in this study, our system’s limitations should be acknowledged. While this study focuses primarily on improving evidence retrieval and frame-semantic parsing, it does not comprehensively evaluate different fact verification approaches. Fact verification is critical for automatic fact-checking, and a deeper investigation into various verification strategies, including fine-tuned models and LLM-based approaches, could further enhance system performance.
Another limitation is that our system’s effectiveness is constrained by the availability of reliable, manually curated data. For claims that lack relevant data in our databases (e.g., OECD tables or U.S. congressional records), the system cannot retrieve evidence, which limits its fact-checking capabilities. This limitation demonstrates the importance of expanding data sources and improving coverage in future iterations of the system.
Ethics and Risks
Automated fact-checking systems, such as the one presented in this work, bring both opportunities and ethical challenges. One key concern is the potential spread of misinformation due to model limitations or errors. Users may over-rely on AI verdicts, leading to the amplification of false positives or negatives, especially in politically sensitive contexts. Biases present in both the models (e.g., GPT-4o-mini, RoBERTa) and the datasets (e.g., PolitiFact, OECD data) could skew fact-checking outcomes, favoring certain political narratives or ideologies.
Another concern is the transparency and accountability of AI systems. If users are unaware of how models arrive at their conclusions, it may be difficult to hold these systems accountable for erroneous outcomes. This opacity could diminish trust in both the fact-checking tool and the institutions that deploy it. Moreover, the collection and processing of political data raise potential privacy concerns, especially regarding the use of public records in ways individuals may not expect.
Marginalized communities may also be disproportionately affected by such systems, as they might misinterpret claims relevant to those groups or lack sufficient representation in training data. Similarly, the system could be exploited by adversaries who craft ambiguous or misleading claims designed to confuse AI, leading to manipulated fact-checks.
Lastly, the over-dependence on specific sources like PolitiFact or congress.gov could limit the tool’s scope. Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
Appendix A Studied Frames
Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings.
Frame Frame Element Frame Definition
Occupy_rank Dimension Items occupying a certain Rank within a hierarchy.
Occupy_rank_via_superlatives Dimension An Item occupying a Rank specified by a superlative.
Comparing_two_entities Comparison_criterion Comparing two entities using a Comparison_criterion and qualifying with a Degree.
Comparing_at_two_different_ points_in_time Comparison_criterion Comparing an Entity with itself at two different points in time using a Comparison_criterion and qualifying with a Degree.
Cause_change_of_position_ on_a_scale Item Words that indicate an Agent or Cause affects the position of an Item on a scale.
Capability Event An Entity meets the pre-conditions for participating in an Event.
Table 6: Frames used for OECD Case Study, along with the Frame Elements extracted from corresponding Frames.
Appendix B Reproducibility
We trained our frame-semantic parser using a NVIDIA GTX 1080 Ti in about 6 hours on a single run without any hyperparameter tuning. Github Copilot was used in the development of some of our code.
B.1 Software License
Our source code and data are released with the MIT License.
Model Parameters URL
msmarco-distilbert-base-tas-b 66M msmarco-distilbert-base-tas-b
msmarco-roberta-base-v2 125M msmarco-roberta-base-v2
msmarco-roberta-base-v3 125M msmarco-roberta-base-v3
stella_en_1.5B_v5 1.5B stella_en_1.5B_v5
text-embedding-3-large unknown text-embedding-3-large
Table 7: Models and Their Corresponding Hugging Face URLs
B.2 Models
Unless stated otherwise, all models used are the “base” variant. All semantic similarity models are available via the sentence-transformers library. [^9] https://huggingface.co/sentence-transformers Table 7 contains all of the models used and their details.
Claim Predicted Verdict Explanation Fact Check
Americans work way more than an average of industrialized countries around the world. Mostly True The data shows that the United States has an annual average of 1829 hours worked per person, which is higher than many other industrialized countries in the dataset, such as Germany (1400.4 hours), France (1518.0 hours), and Japan (1729.0 hours). However, there are some countries like Mexico (2229.0 hours), Korea (2076.0 hours), and Chile (1994.0 hours) with higher working hours than the U.S. This indicates that while Americans work more than many other industrialized countries, they do not work more than all. The claim is mostly true but lacks the nuance that there are industrialized countries where people work more hours. PolitiFact
The United States is the ONLY industrialized country without universal healthcare. True The retrieved data indicates that countries such as Italy, the Netherlands, Spain, Canada, Korea, Australia, France, Germany, the United Kingdom, and Sweden all have near-universal health insurance coverage through government or compulsory schemes, with coverage rates often very close to 100% of the population. In contrast, the United States has only 38.1% of its population covered by government or compulsory health insurance, suggesting that it lacks universal healthcare coverage. This supports the claim that the United States is the only industrialized nation without such a system based on the data provided. PolitiFact
Table 8: OECD fact check example predictions and explanations
Appendix C Examples
In this section, we provide a few example outputs of our system to give a better understanding of the outputs of our models.
Claim Bill Title Vote Alignment Prediction
Marco Rubio voted against the bipartisan Violence Against Women Act 112 S 1925 Violence Against Women Reauthorization Act of 2012 Nay Supports True
17 HR 3233 National Commission to Investigate the January 6 Attack on the United States Capitol Complex Act Nay Irrelevant
117 HR 350 Domestic Terrorism Prevention Act of 2022 Nay Irrelevant
Chuck Grassley was voting to slash Medicare when voting against the debt ceiling bill 117 S 610 Protecting Medicare and American Farmers from Sequester Cuts Act Nay Supports Mostly False
117 S 1301 Promoting Physical Activity for Americans Act Nay Irrelevant
17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
Table 9: Examples of voting-related claims with the corresponding retrieved bills, votes on the bill, the vote-claim alignment, and fact verification prediction.
C.1 Vote-Claim Alignment
We include two sample claims and their corresponding top three retrieved bills along with the vote-claim alignment for each bill in Table 9.
C.2 Fact Verification Predictions
Vote
We include two claims and their corresponding predictions based on the retrieved bills in Table 9.
OECD
We include two predictions and their explanations on OECD claims in Table 8.
Appendix D LLM Prompts
D.1 Frame-Semantic Parsing
Listing 1: Frame-semantic parser for the Vote frame.
Identify if a Vote semantic frame is evoked in a given sentence. If it is, extract and list the relevant frame elements associated with the voting event.
A Vote frame is defined as: An Agent makes a voting decision on an Issue.
Frame elements to identify:
- Agent: The conscious entity (usually a person) executing the voting decision.
- Issue: The subject or matter that the Agent is voting on with a particular position.
- Side: Additional person(s) involved in the voting on the same Issue alongside the Agent.
- Position: The stance the Agent takes, expressing whether their vote is in favor or against the Issue.
- Frequency: How often the Agent has made this voting decision regarding the Issue.
- Time: When the Agent performed the voting decision.
- Place: Location where the vote took place.
- Support_rate: The percentage of the Agents total votes aligning with the Side over a set period.
# Notes
- If a frame element is not mentioned in the sentence, use "" for its value.
- Carefully distinguish between elements; some may have overlapping characteristics.
- Consider synonyms or variations of terms related to voting when analyzing the sentence.
- Frame elements should quote exactly from the input
Listing 2: Frame-semantic parser for OECD frames.
Identify semantic frames evoked by a given factual claim and extract relevant frame elements for each identified frame.
Consider the predefined semantic frames and their elements:
- Occupy_rank_via_superlatives:
- Item: Entity occupying the rank.
- Rank: Rank held, often defined by a superlative.
- Dimension: Aspect along which the ranking occurs (e.g., speed, age).
- Comparison_set: Group of entities being compared.
- Time: Time period when the item occupies the rank.
- Occupy_rank:
- Item: Entity occupying the rank.
- Rank: Rank held.
- Dimension: Criterion used for ranking.
- Comparison_set: Set of entities being ranked.
- Time: Time period when the item holds the rank.
- Change_position_on_a_scale:
- Item: Entity whose position on a scale changes.
- Attribute: Property or scale of the change.
- Difference: Amount of change.
- Final_value: Final position on the scale.
- Initial_value: Initial position on the scale.
- Path: Progression between positions on the scale.
- Speed: Rate of change.
- Correlated_variable: Related variable that changes with the Attribute.
- Manner: Method of performing the action.
- Degree: Extent to which the change occurs.
- Circumstances: Context or conditions for the change.
- Result: Outcome of the change.
- Group: Collection of Items undergoing change.
- Time: Time frame during which the change occurs.
- Duration: Length of time for the change to happen.
- Value_range: Range of values along the scale.
- Initial_correlate: State corresponding to the Initial_value.
- Final_correlate: State corresponding to the Final_value.
- Place: Location where the Attribute is measured.
- Initial_state: State of the Item before the change.
- Final_state: State of the Item after the change.
- Period_of_iterations: Duration from start to stop of repeated changes.
- Particular_iteration: Specific instance of a repeated event.
- Containing_event: Broader event encompassing the change.
- Explanation: Rationale for the change.
- Comparing_two_entities:
- Entity_1: First entity in comparison.
- Entity_2: Entity serving as the comparison point.
- Comparison_criterion: Criterion for comparison.
- Degree: Extent of the comparison.
- Time: Period over which the comparison takes place.
- Comparing_at_two_different_ points_in_time:
- Entity: Entity compared to itself at different times.
- First_point_in_time: First time period in comparison.
- Second_point_in_time: Second time period in comparison.
- Comparison_criterion: Criterion being compared.
- Degree: Extent of the difference.
- Difference: Magnitude of change between the two times.
# Steps
1. Carefully read and analyze the factual claim.
2. Determine which semantic frame(s) are invoked by the claim.
3. For each evoked frame, extract the appropriate frame elements directly from the claim.
4. Document each evoked frame and the extracted elements.
# Output Format
- List each evoked frame followed by its extracted elements in a structured form.
- Ensure all extracted frame elements are clearly labeled and match the exact inputs.
# Notes
- Some claims may trigger multiple frames; ensure all applicable frames are evaluated.
- If specific frame elements cannot be determined, note their absence for completeness.
- If a frame element does not explicitly appear in the claim, leave it blank.
D.2 Claim-Bill Alignment
Listing 3: Claim-Bill Alignment
Given the following factual claim, bill summary, and vote on the bill, evaluate whether the content of the bill summary and the voting record align with the given claim. You may consider factors such as the main objectives of the bill and unintended or implicit consequences. Your task is to determine if the information provided in the bill summary and the voting record supports or refutes the given factual claim. Return your explanation and one of the following labels in JSON format.
Bill Summary:
{summary}
Vote: {vote_type}
Claim: {claim}
Labels:
Supports - The vote on this bill directly or indirectly supports the claim.
Refutes - The vote on this bill explicitly refutes the claim.
Inconclusive - The vote on this bill does not provide enough information to support or refute the claim.
Irrelevant - The vote on this bill is not relevant to the claim at all.
D.3 OECD Data Query
Listing 4: Retrieve Data for Fact-Checking Claim
Your task is to write a Python function named retrieve_data() that retrieves data for fact-checking the claim:
{claim}
Given the following database schemas from OECD_Data.db:
{table_descriptions}
Instructions:
- Use LIKE only for textual columns. For numerical columns (e.g., year, value), use appropriate comparison operators like = or <=.
- Do not modify the database.
- The function must contain all necessary imports inside it and take no parameters, though passing parameters (e.g., file paths or claim details) should be considered in future iterations.
- Return the data in pandas DataFrames. Multiple DataFrames can be returned as a list if needed.
- If no relevant data is found, return Data is not Available’.
- Use only actual values found in the columns. If aggregation is possible (e.g., summing or averaging categories), return aggregated results, unless the claim specifies otherwise. Use appropriate aggregation methods based on the data.
- Always return relevant columns (avoid ’*’ when possible). Select the necessary columns based on the claim.
- Treat we in the claim as referring to United States’, but consider the context of the claim for potential exceptions.
- If any unit_of_measure is National currency’, use the USD_value column instead of value’.
- Use the nearest available date based on the provided dictionary: {nearest_dates}. If no close date is found, return Data is not Available’.
- For multiple tables, create multiple queries as needed. If combining data from multiple tables is required, ensure consistency in merging and handling differences in structure.
- Do not filter by country or unit_of_measure’; that will be handled later.
Output:
- Return a list of pandas DataFrames or Data is not Available if no relevant data is found.
Listing 5: Clean Data for Fact-Checking Claim
Your task is to write a Python code function named clean_data()‘ to filter and clean data for fact-checking the claim {claim} from OECD_Data.db‘.
Extracted dataframes from the previous step:
{list_of_dfs}
These dataframes were generated using this code:
{code}
The schema for these dataframes is:
{schema_str}
Notes:
- The function must have all imports inside it and take no parameters.
- This function will run independently of the given code, meaning you need to re-extract the data from the database to clean it for fact-checking the claim:
{claim}
- Do not modify the original way the data was extracted unless the input code is very incorrect; just add more filters/conditions based on the claim.
- Keep the same aggregated structure from the original extraction code unless the claim specifies otherwise.
- Filter by the nearest date: {nearest_date}.
- Exclude tables only if they are irrelevant to the claim.
- Use only the actual tables and data that were extracted; do not make up new tables.
- Add more filters to narrow down to claim-relevant metrics; try to have one value per non-numeric column wherever possible.
- Claim-relevant metrics may appear in columns other than ‘’measure’‘, so reference the claim to identify the appropriate columns to filter by.
- Use the retrieval code to identify useful information for your filters, if needed.
- When filtering by ‘’unit_of_measure’‘, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
Output:
- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return ‘’Data is not Available’‘.
D.4 Fact Verification
Listing 6: Vote - Verify Claim and Provide Verdict
You are given the following claim, and 5 bills along with the bill title, a short summary of the bill, the vote cast on the bill and the alignment of the bill with the claim.
Alignment here means whether an individual bill supports, refines, or is irrelevant to the claim.
Your task is to give one of the given fact-check labels to the claim based on the evidence which are the bills.
You may consider factors such as the main objectives of the bill and unintended or implicit consequences.
Return your explanation and one of the following labels in JSON format.
Claim: {claim}
Bill Title 1: {bill_title_1}
Bill Summary 1: {bill_summary_1}
Vote 1: {vote_type_1}
Alignment1: {alignment_1}
Bill Title N: {bill_title_N}
Bill Summary N: {bill_summary_N}
Vote N: {vote_type_N}
Alignment N: {alignment_N}
Labels:
True: The given bills support the claim.
MostlyTrue: The given bills mostly support the claim.
HalfTrue: The given bills can only partly support or refute claim.
MostlyFalse: The given bills mostly refute the claim.
False: The given bills refute the claim.
Irrelevant: The given bills are not relevant to the claim.
Return the JSON object with the label and the explanation. The fields should be Label and Explanation’.
Listing 7: OECD - Verify Claim and Provide Verdict
Your task is to verify the claim using the retrieved data and provide a verdict.
Claim: {claim}
Retrieved Data:
{formatted_data}
Instructions:
- Use only the provided data, regardless of its perceived relevance, to analyze the claim.
- The verdict must be based solely on the retrieved data. Do not rely on external knowledge.
- Ensure you consider the spirit of the claim in the fact-check and not just the precise verbage.
Verdict Categories:
- True: The statement is fully accurate, with no significant information missing.
- Mostly True: The statement is accurate but requires clarification or additional context.
- Half-True: The statement is partially accurate but omits important details or misrepresents the context.
- Mostly False: The statement contains some truth but overlooks key facts that would significantly alter the impression.
- False: The statement is completely inaccurate.
Output:
- Provide a verdict in the format:
Verdict: [False, Mostly False, Half-True, Mostly True, True]; Explanation: [Your reasoning].